OpenStudy (anonymous):

. Which of the following is the equation of a circle with a radius of 2 and its center at (4, 3)? (Points : 5) (x 4)2 + (y + 3)2 = 2 (x + 4)2 + (y 3)2 = 2 (x 4)2 + (y + 3)2 = 4 (x + 4)2 + (y 3)2 = 4 I think it is b but i am unsure

OpenStudy (anonymous):

is it (x - 3)^2 + (y - 2)^2 = 4

OpenStudy (anonymous):

i mean -4

jimthompson5910 (jim_thompson5910):

the center is (4, 3), so the left side should be (x - 4)^2 + (y - 3)^2

jimthompson5910 (jim_thompson5910):

idk how you got (x - 3)^2 + (y - 2)^2

OpenStudy (anonymous):

lol yeah I kinda have a messed up calculator :( lol

jimthompson5910 (jim_thompson5910):

no need for a calculator since you're just plugging in (or replacing) variables with numbers

OpenStudy (anonymous):

thats probaby how im confusing myself ugh lol

jimthompson5910 (jim_thompson5910):

maybe, i guess that's a sign you're probably too dependent on your calculator, but idk

jimthompson5910 (jim_thompson5910):

but all you're doing is going from (x-h)^2+(y-k)^2=r^2 to (x-4)^2+(y-3)^2=2^2 .. replace h with 4, k with 3, r with 2 (x-4)^2+(y-3)^2=4

OpenStudy (anonymous):

yeah i kinda am lol ever since highschool all the teachers would say is use your calculator not ur head you could be wrong but the calculata is always right lol

jimthompson5910 (jim_thompson5910):

well if you type it in wrong, then the calculator is wrong it does exactly what it's told to do

OpenStudy (anonymous):

they also say if you dont know the answer then "guess" but i feel like thats kinda stupid lol

OpenStudy (anonymous):

yeah so true

jimthompson5910 (jim_thompson5910):

and calculators don't handle everything...otherwise, why not just hand the job over to the calculator

jimthompson5910 (jim_thompson5910):

well they're teaching to the test because you can't guess if it isn't multiple choice